# Aouad, Ji, and Shaposhnik

### v1

@@ 5.4 When is P-opening worthwhile?
We next conduct experiments that shed light on the structure of optimal policies and further characterize when P-opening is worthwhile.
To this end, we define the notion of P-ratio as the average percentage of P-openings conducted by the optimal policy of all opening actions. A higher P-ratio implies that the optimal policy applies P-opening more frequently. We compute the P-ratio for a set of problem instances and plot it against various problem characteristics. For brevity, we describe the main results below and postpone the complete details of the experiment and the figures to Appendix 11.1.
One of our main findings is that greater homogeneity of boxes in terms of their opening thresholds (i.e., boxes $i$ and $j$ are “homogeneous” when the tuples $(\sigma^{F}_{i},\sigma^{P}_{i})$ and $(\sigma^{F}_{j},\sigma^{P}_{j})$ are close) is correlated with more frequent partial opening. This aligns with intuition, since when boxes are very different, their relative value can be determined without partial inspection; hence, F-opening can simply be conducted as a necessity prior to selection.
Additionally, in line with our theoretical results, we observe an increased tendency to partially open when there is a higher proportion of boxes whose P-thresholds exceed their corresponding F-thresholds. This finding underscores the role of opening thresholds in capturing the value of corresponding inspection modes.
Finally, in the case of identical boxes (homogeneity is maximal), the tendency to partially open increases in the number of boxes. This suggests that when the pool of alternatives is larger, partial opening becomes more significant—assuming partial opening is worthwhile to begin with.
5.5 Evaluation of generalized index policies
Section 5.3 compares different policies and shows that $\pi^{F^{*},P^{*}}$ and $\pi^{index}$ stand out in their near-optimal performance. These are index policies that assign a static value to each box-action, and for every system state, select the box and action whose index value is maximal. Although the theoretical analysis showed that index policies need not be optimal, the numerical experiments suggest that in practice they perform optimally in most cases. This raises a question about the extent to which index policies, or their generalizations, can accurately represent optimal policies. In what follows, we describe the results of a numerical study in which we fit index policies to numerically obtained optimal solution and evaluate the quality of the fit (see Appendix 11.2 for the details).
First, we observe that, for most problem instances, there is no index rule that perfectly fits the structure of the optimal policy. Contrasting this finding with Table 4 suggests that, in many cases, index policies can prescribe optimal decisions that follow relatively few narrow paths throughout the state space that begin with the initial state and end at stopping. However, if for some reason a deviation from the paths occurs, then the index policies may prescribe suboptimal decisions. This is in contrast to the standard Pandora’s box problem in which the optimal index policy remains optimal for every system state.
Second, we show that a more general policy that consists of multiple index rules better represents the optimal policy. This policy is defined using multiple index rules and a classifier which determines which index rule to follow based on the current system state. Specifically, we evaluate policies that are defined using two index rules and a linear classifier that applies a linear function to a numeric encoding of the system state to determine which index rule to follow. Tables 6 and 7 in Appendix 11.2 show that our generalized index policies offer a much better fit to the optimal policy. These policies can therefore offer an interpretable representation of optimal policies (as alternating between two index rules), as well as a smaller set of policies to consider for optimization.
@@ 6 Conclusions and Future Directions
The work studies a generalized Pandora’s box problem that captures a tradeoff related to acquiring accurate and costly versus inaccurate but cheaper information about the value of competing alternatives. We proved that the problem is NP-hard, but nevertheless, using a synthesis of techniques, we identified structural properties of the optimal policy and provided insights into the key drivers of optimal information acquisition decisions. We showed that intuitive threshold-based policies that extend the Pandora’s box solution describe optimal decisions and form the basis of approximate solutions. Extensive numerical experiments show that this characterization almost entirely explains the optimal policy, and that the approximate policies perform extremely well.
As our work focused on the simplest yet sufficiently rich context where the aforementioned tradeoff occurs, it would be interesting to consider more general settings. For example, a richer information model that allows multiple types of repeated inspections, inspection costs that decrease after inspections take place, and the option to select partially opened boxes, all could be interesting directions for future work.

### v2

@@ 5.4 When is P-opening worthwhile?
We next conduct experiments that shed light on the structure of optimal policies and further characterize when P-opening is worthwhile. To this end, we define the notion of P-ratio as the average percentage of P-openings conducted by the optimal policy of all opening actions. A higher P-ratio implies that the optimal policy applies P-opening more frequently. We compute the P-ratio for a set of problem instances and plot it against various problem characteristics. For brevity, we describe the main results below and postpone the complete details of the experiment and the figures to Appendix 12. One of our main findings is that greater homogeneity of boxes in terms of their opening thresholds (i.e., boxes $i$ and $j$ are “homogeneous” when the tuples $(\sigma^{F}_{i},\sigma^{P}_{i})$ and $(\sigma^{F}_{j},\sigma^{P}_{j})$ are close) is correlated with more frequent partial opening. This aligns with intuition, since when boxes are very different, their relative value can be determined without partial inspection; hence, F-opening can simply be conducted as a necessity prior to selection. Additionally, in line with our theoretical results, we observe an increased tendency to partially open when there is a higher proportion of boxes whose P-thresholds exceed their corresponding F-thresholds. This finding underscores the role of opening thresholds in capturing the value of corresponding inspection modes. Finally, in the case of identical boxes (homogeneity is maximal), the tendency to partially open increases in the number of boxes. This suggests that when the pool of alternatives is larger, partial opening becomes more significant—assuming partial opening is worthwhile to begin with.